# Approximating the coefficients of the Bessel functions

### v4

@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $\mu_{1}$, $\mu_{2}$, and $\mu$ are Borel probability measures over $\mathbb{R}^{N}$ such that $G_{\mu}^{\mathcal{R}(\theta)}=G_{\mu_{1}}^{\mathcal{R}(\theta)}G_{\mu_{2}}^{\mathcal{R}(\theta)}$ over an open subset of $i\mathbb{R}^{N}$. Then, every element of $\text{supp}(\mu)$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$ for $i\in\{1,2\}$.
@@
For the sake of contradiction, assume that the distance between $u$ and $\text{supp}(\mu_{a_{i}}^{\text{sym}})$ for any element $a_{i}\in\text{supp}(\mu_{i})$ is greater than $2r$. Then, for all $a_{i}\in\text{supp}(\mu_{i})$, $\Pr_{\epsilon_{i}\sim\mu_{a_{i}}^{\text{sym}}}[\epsilon_{i}\in B(u,r)]=0$, which is a contradiction to (34). Recall from Theorem˜1.4 that for $a\in\mathbb{R}^{N}$, the support of $\mu_{a}^{\text{sym}}$ is a subset of the convex hull of $H(\mathcal{R})a$. By considering when $r\rightarrow 0$, we have that $u$ must be in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$. ∎
@@
      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
@@
      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
    For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.
@@
    Suppose $\theta\in\mathbb{C}$ is positive. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true. For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in distribution and in terms of moments.
@@
(A): We can follow the proof of Corollary˜1.7 that involves first proving convergence in terms of moments and using this fact to justify convergence in distribution. Instead of using Theorem˜1.1, we use Corollary˜10.17.
(B): We use Lemma˜10.21 to deduce that for $N\geq 1$, $\mu^{N}$ is compactly supported and therefore exponentially decaying. Afterwards, we follow the proof of (A) for the remaining steps of the proof. ∎
@@
      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
@@
      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
@@
    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
@@
    Suppose $\theta_{0}$ and $\theta_{1}$ are nonnegative. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true. For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
@@
      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
@@
      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
    For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
@@
    Suppose $\theta\in\mathbb{C}$ is positive. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true. For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
@@
As an example, we consider the Dyson Brownian motion with a random initial condition. For $\theta>0$, the Dyson Brownian motion $DBM(\theta)$ over $\mathbb{R}^{N}$ with initial values $(a_{1}(0),\ldots,a_{N}(0))$ is the unique strong solution to
@@
For $t\geq 0$, let $\mu_{t}(a)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ with initial values $a\in\mathbb{R}^{N}$. From [airy_beta]*Lemma 3.8,
@@
Suppose $\eta$ is a probability measure over $\mathbb{R}^{N}$. Let $\mu_{t}(\eta)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ if we assume that the initial values $(a_{1}(0),\ldots,a_{N}(0))$ are distributed according to $\eta$ independently of the Brownian motions. In particular, we define
$\mu_{t}(\eta)\triangleq\operatorname{\mathbb{E}}_{a\sim\eta}[\mu_{t}(a)].$
@@
Assume that $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$. Then, for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$ and
$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x)$
@@
First, we show that for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$. This is obviously true if $t=0$, so assume that $t>0$.
@@
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2t}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
@@
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
This implies that $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$.
For the computation of the Bessel generating function $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}$, we have that
$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{2N}}J_{a}^{A^{N-1}(\theta)}(x)d\mu_{t}(\eta)(a).$
for $x\in\mathbb{C}^{N}$. Because $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$, we can use part (e) of Lemma˜2.6 and Fubini’s theorem to deduce that for $x\in B_{R}(0)$,
$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{N}}\left(\int_{\mathbb{R}^{N}}J_{a(t)}^{A^{N-1}(\theta)}(x)d\mu_{t}(a)(a(t))\right)d\eta(a).$
Afterwards, applying (36) and part (A) of Lemma˜10.15 implies that this expression equals
@@
over $\mathbb{C}^{N}$.
@@
Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.27 implies that
(37) $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)G^{A^{N-1}(\theta)}_{H_{\theta}^{\sqrt{\theta t}}}(x)$
over a neighborhood of the origin for exponentially decaying Borel probability measures $\eta$. This example satisfies Conjecture˜1.6 for randomly distributed $a_{1}$ and $a_{2}$.
@@
Suppose $\theta>0$ and that for $N\geq 1$, $\eta_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{R}^{N}$ such that as $N\rightarrow\infty$, $\operatorname{\mathbb{E}}_{a\sim\eta_{N}}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges in terms of moments to the distribution $\eta$ over $\mathbb{R}$. Then, for $\alpha\geq 0$, $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges in terms of moments to the free convolution of $\eta$ and the semicircular law multiplied by $\sqrt{\frac{\alpha}{\theta}}$ as $N\rightarrow\infty$. If $\eta$ is compactly supported, then $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges weakly to the free convolution as $N\rightarrow\infty$.
@@
Assume that $\alpha>0$. We can follow the method of the proof of Corollary˜10.22. First, since condition (b) of Corollary˜10.17 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well. The next step is to deduce that $\mu_{\alpha N}(\eta_{N})$ is exponentially decaying for all $N\geq 1$. This does not follow from (37) and Lemma˜10.21 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
@@
Assume that $\eta_{N}$ exponentially decays at rate $R$. By Lemma˜10.27, $\mu_{\alpha N}(\eta_{N})$ exponentially decays at any rate less than $R$. Then, because
$G_{\mu_{\alpha N}(\eta_{N})}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta_{N}}^{A^{N-1}(\theta)}(x)$

### v5

@@ 10.5. Bessel generating functions for exponentially decaying probability measures
Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $\mu_{1}$, $\mu_{2}$, and $\mu$ are exponentially decaying Borel probability measures over $\mathbb{R}^{N}$ such that $G_{\mu}^{\mathcal{R}(\theta)}=G_{\mu_{1}}^{\mathcal{R}(\theta)}G_{\mu_{2}}^{\mathcal{R}(\theta)}$ over an open subset of $i\mathbb{R}^{N}$. Then, every element of $\text{supp}(\mu)$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$ for $i\in\{1,2\}$.
@@
For the sake of contradiction, assume that the distance between $u$ and $\text{supp}(\mu_{a_{i}}^{\text{sym}})$ for any element $a_{i}\in\text{supp}(\mu_{i})$ is greater than $2r$. Then, for all $a_{i}\in\text{supp}(\mu_{i})$, $\Pr_{\epsilon_{i}\sim\mu_{a_{i}}^{\text{sym}}}[\epsilon_{i}\in B(u,r)]=0$, which is a contradiction to (25). Recall from Theorem˜1.3 that for $a\in\mathbb{R}^{N}$, the support of $\mu_{a}^{\text{sym}}$ is a subset of the convex hull of $H(\mathcal{R})a$. By considering when $r\rightarrow 0$, we have that $u$ must be in the convex hull of $H(\mathcal{R})a_{i}$ for some $a_{i}\in\text{supp}(\mu_{i})$. ∎
@@
      The measure $\mu_{a}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{a}$ over $\mathbb{C}$ with finite moments in terms of power sums.
@@
      The measure $\mu_{b}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{b}$ over $\mathbb{C}$ with finite moments in terms of power sums.
    For $N\geq 2$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mu_{N}(x\mapsto\frac{x}{\theta N})$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ with finite moments in terms of power sums.
@@
    Suppose $\theta\in\mathbb{C}$ is positive. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 2$ and $\mu_{a}$ and $\mu_{b}$ are Borel probability measures over $\mathbb{R}$ with finite moments such that conditions 1 and 2 from (A) are true. For $N\geq 2$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{A^{N-1}(\theta)}=G_{\mu_{a}^{N}}^{A^{N-1}(\theta)}G_{\mu_{b}^{N}}^{A^{N-1}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mu_{N}(x\mapsto\frac{x}{\theta N})$ converges to the free convolution of $\mu_{a}$ and $\mu_{b}$ in terms of power sums.
@@
(A): We can follow the proof of Corollary˜1.7. Instead of using Theorem˜1.2, we use Corollary˜10.18.
(B): We use Lemma˜10.23 to deduce that for $N\geq 2$, $\mu^{N}$ is compactly supported and therefore exponentially decaying. Afterwards, we follow the proof of (A). ∎
@@
      The measure $\mu_{a}^{N}(x\mapsto\frac{x}{\theta_{0}N})$ converges to the Borel probability measure $\mu_{a}$ over $\mathbb{C}$ with finite moments in terms of power sums.
@@
      The measure $\mu_{b}^{N}(x\mapsto\frac{x}{\theta_{0}N})$ converges to the Borel probability measure $\mu_{b}$ over $\mathbb{C}$ with finite moments in terms of power sums.
@@
    Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta_{0}N})$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in terms of power sums.
@@
    Suppose $\theta_{0}$ and $\theta_{1}$ are nonnegative. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 2$ and $\mu_{a}$ and $\mu_{b}$ are Borel probability measures over $\mathbb{R}$ with finite moments such that conditions 1 and 2 from (A) are true. For $N\geq 2$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
    Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta_{0}N})$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in terms of power sums.
@@
      The measure $\mu_{a}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{a}$ over $\mathbb{C}$ with finite moments in terms of power sums.
@@
      The measure $\mu_{b}^{N}(x\mapsto\frac{x}{\theta N})$ converges to the Borel probability measure $\mu_{b}$ over $\mathbb{C}$ with finite moments in terms of power sums.
    For $N\geq 2$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin. Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta N})$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in terms of power sums.
@@
    Suppose $\theta\in\mathbb{C}$ is positive. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 2$ and $\mu_{a}$ and $\mu_{b}$ are Borel probability measures over $\mathbb{R}$ with finite moments such that conditions 1 and 2 from (A) are true. For $N\geq 2$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin. Then, the measure $\frac{1}{2}\mu^{N}(x\mapsto\frac{x}{\theta N})+\frac{1}{2}\mu^{N}(x\mapsto-\frac{x}{\theta N})$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in terms of power sums.
@@
As an example, we consider the Dyson Brownian motion with a random initial condition. For $\theta\geq\frac{1}{2}$, the Dyson Brownian motion $DBM(\theta)$ over $\mathbb{R}^{N}$ with initial values $(a_{1}(0),\ldots,a_{N}(0))$ is the unique strong solution to
@@
For $t\geq 0$, let $\mu^{\theta}_{t}(a)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ with initial values $a\in\mathbb{R}^{N}$. From [airy_beta]*Lemma 3.8,
@@
Suppose $\eta$ is a probability measure over $\mathbb{R}^{N}$. Let $\mu^{\theta}_{t}(\eta)$ denote the distribution of $(a_{1}(t),\ldots,a_{N}(t))$ if we assume that the initial values $(a_{1}(0),\ldots,a_{N}(0))$ are distributed according to $\eta$ independently of the Brownian motions. In particular, we define
$\mu^{\theta}_{t}(\eta)\triangleq\operatorname{\mathbb{E}}_{a\sim\eta}[\mu^{\theta}_{t}(a)].$
@@
Assume that $\theta\geq\frac{1}{2}$ and $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$. Then, for all $t\geq 0$, $\mu^{\theta}_{t}(\eta)$ exponentially decays at any rate less than $R$ and
$G_{\mu^{\theta}_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x)$
@@
First, we show that for all $t\geq 0$, $\mu^{\theta}_{t}(\eta)$ exponentially decays at any rate less than $R$. This is obviously true if $t=0$, so assume that $t>0$.
@@
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2t}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
@@
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}^{\theta}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
This implies that $\mu_{t}^{\theta}(\eta)$ exponentially decays at any rate less than $R$.
For the computation of the Bessel generating function $G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}$, we have that
$G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{2N}}J_{a}^{A^{N-1}(\theta)}(x)d\mu_{t}^{\theta}(\eta)(a).$
for $x\in\mathbb{C}^{N}$. Because $\mu_{t}^{\theta}(\eta)$ exponentially decays at any rate less than $R$, we can use part (e) of Lemma˜2.6 and Fubini’s theorem to deduce that for $x\in B_{R}(0)$,
$G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{N}}\left(\int_{\mathbb{R}^{N}}J_{a(t)}^{A^{N-1}(\theta)}(x)d\mu_{t}^{\theta}(a)(a(t))\right)d\eta(a).$
Afterwards, applying (27) and part (A) of Lemma˜10.16 implies that this expression equals
@@
over $\mathbb{C}^{N}$; this formula also follows from Theorem˜8.1.
@@
Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.29 implies that
(28) $G_{\mu_{t}^{\theta}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)G^{A^{N-1}(\theta)}_{H_{\theta}^{\sqrt{\theta t}}}(x)$
over a neighborhood of the origin for exponentially decaying Borel probability measures $\eta$. This example satisfies Conjecture˜1.5 for randomly distributed $a_{1}$ and $a_{2}$.
@@
Suppose $\theta\geq\frac{1}{2}$ and that for $N\geq 2$, $\eta_{N}$ is an exponentially decaying Borel probability measure over $\mathbb{R}^{N}$ such that $\eta_{N}(x\mapsto\frac{x}{\theta N})$ converges in terms of power sums to the distribution $\eta$ over $\mathbb{R}$ with finite moments. Then, for $\alpha\geq 0$, $\mu_{\alpha N}^{\theta}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ converges in terms of power sums to the free convolution of $\eta$ and the semicircle law multiplied by $\sqrt{\frac{\alpha}{\theta}}$ as $N\rightarrow\infty$.
@@
Assume that $\alpha>0$. We can follow the method of the proof of Corollary˜10.24. First, since condition (b) of Corollary˜10.18 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well. The next step is to deduce that $\mu_{\alpha N}^{\theta}(\eta_{N})$ is exponentially decaying for all $N\geq 2$. This does not follow from (28) and Lemma˜10.23 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
@@
Assume that $\eta_{N}$ exponentially decays at rate $R$. By Lemma˜10.29, $\mu^{\theta}_{\alpha N}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ exponentially decays at any rate less than $R$. Then, because
$G_{\mu_{\alpha N}^{\theta}(\eta_{N})}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta_{N}}^{A^{N-1}(\theta)}(x)$